OpenStudy (anonymous):

a) Calculate the following (geometric) sums and series, respectively. ii) \(\Large \sum_{n=2}^{\infty}3^{1-n}\)

OpenStudy (anonymous):

a=1/3 q=1/3

OpenStudy (anonymous):

can you do the complete solution pls, i have also some solution but i must be sure its correct, tomorrow i must give it up by mentor and its important

OpenStudy (anonymous):

ok \[\huge \sum_{n=2}^{\infty} 3^{1-n}=\frac{1}{3}+\frac{1}{3^2}+\frac{1}{3^3}+... \]

OpenStudy (anonymous):

yes i have the same one, can i give like that to by mentor ?

OpenStudy (anonymous):

am i right?

OpenStudy (anonymous):

what you think is it possible to give like that to the mentor ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

what we need to add more ?

OpenStudy (anonymous):

for geometric series like \[\large a+aq+aq^2+aq^3+...\] we have \[\huge s=\frac{a}{1-q}\]

OpenStudy (anonymous):

hmm ...

OpenStudy (anonymous):

ook

OpenStudy (anonymous):

u just recognize a and q

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

welcome